Benxoic acid, C7H5O2H is a weak monoprotic acid (Ka = 6.3 * 10-5). Consider a titration between 20.0 mL of 0.100M benzoic acid solution with 0.200 M sodium hydroxide, NaOH.

a.) What volume of NaOH is required to reach the equivalence point?

b.) Calculate the pH of the solution at equivalence point
Answer #1

cy H, H: molarity = 0.1M | ka= 6-3 XL05 . volume arom1=0.02L Loka_-log ka = -log (6-3) pkaa 4.2 1 pkW 214 719=-109 (6-38105)pH formulas for salts weak acid + strong base ->salft iho pH = Ź [pew tpka + loge] strongacid & weakbase salf to a pH = ₃ [